# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v1

@@ 5 Finite size objects at zero pressure
An object with size $R\gg 1/m_{a}$ is stable at almost zero pressure when
@@
$\begin{split}M&\simeq\frac{4\pi R^{3}}{3}\frac{m_{n}(5m_{n}\varepsilon m_{\pi}^{2}f_{\pi}^{2})^{3/5}}{(3\pi^{2})^{2/5}}\left[1-f(\pi)+\frac{m_{u}m_{d}}{2Rm_{a}(m_{u}+m_{d})^{2}}\right]^{3/5}\,,\\ &\simeq 2\times 10^{18}\varepsilon^{3/5}\left(\frac{R}{1m}\right)^{3}\,{\rm kg}\,,\end{split}$ (34)
@@
Figure 11 shows the minimum stable radius for $\pi$-balls made of SNM and PNM and Fig. 12 shows the minimum stable mass, including nuclear interactions from the RMFT model and stabilized by axions. The horizontal dotted lines in Fig. 11 show when $m_{a}R=\pi$ and the domain wall approximation fails. The black dotted lines in Fig. 12 show the mass at which gravity cannot be neglected from Eq. (35). Note that when the constraints on axion parameter space from neutron star measurements we will derive in section 6 is taken into account, these $\pi$-balls should typically have masses larger than the Earth. They will have masses comparable to heavy planets or brown dwarfs, and sizes as large as a few kilometers. Observation of one of these objects would be quite remarkable.
6 Observable implications and constraints
In this section, we estimate constraints on the exceptionally light axion parameter space from neutron star observations. We focus on three observables based on modifications to the crust of an axion-condensed neutron star: thermal relaxation times of the crust in accreting systems (Sec. 6.1), cooling rates of isolated neutron stars based on a diminished envelope, first considered in Ref. nscooling_qcdaxion2024 , (Sec. 6.2), and fractional moment inertia of the crust as inferred from pulsar glitches (Sec. 6.3). All of these rely only on the understanding of neutron star matter and dynamics below saturation density. In Sec. 6.4, we summarize the constraints in the context of the $Z_{\mathcal{N}}$ model of exceptionally light axions.
We note also that the mass-radius relation of axion condensed stars has some unusual features (see Sec. 4.1) which could lead to constraints on axion models. Current measurements of the masses and radii of neutron stars are not yet precise enough to draw conclusions. Next-generation experiments may significantly reduce the error bars and yield a useful constraint if the EOS for a range of densities, and for both $\theta=0$ and $\theta=\pi$, is better understood.
To understand the constraint from observation, we apply the methods discussed in Sec. 4 and specifically the understanding of the axion domain wall established in Sec. 4.2. There are three distinct regions for this calculation. First, when $\theta=\pi$ is the favored phase all the way to zero pressure, the domain wall sits at the surface and forms a thin shell surrounding the rest of the star at negative bulk pressure, resulting in a thinner crust if $\varepsilon$ is not too small (cf. Fig. 10). Second, when the transition from $\theta=\pi$ to $\theta=0$ occurs in the crust of the normal phase, the domain wall connects a homogeneous axion condensed phase to a heterogeneous normal phase and the size of the crust is diminished. Third, when the phase transition to $\theta=\pi$ occurs at a density above the crust-core phase transition, the entire normal crust is present and we draw no constraint.
@@ 6 Observable implications and constraints
DW in core at $n<n_{\rm sat}$ $0.380.27<\varepsilon<$ 0.38 $0.750.56<\varepsilon<$ 0.75 $10.96<\varepsilon<$ 1
6.1 Constraint from crust thermal relaxation in x-ray binaries
Neutron stars in transiently accreting x-ray binaries go through periods of outburst and quiescence, reaching a quasi-steady state on astronomical timescales Brown_1998 . During outburst, matter is rapidly accreted onto the neutron star surface from a companion and pycnonuclear reactions deep in the crust heat the star. On timescales of tens to hundreds of days (see, e.g., Ref. Cackett_2008 ) the crust cools and comes back into thermal equilibrium with the core. Properties of the crust of neutron stars in these x-ray binaries can be inferred by observing the timescale for the crust to thermally relax following an outburst. The thermal timescale is roughly given by $\tau_{\rm th}\simeq C_{V}(\Delta r)^{2}/\kappa$ where $C_{V}$ is the specific heat of the crust, $\kappa$ is the thermal conductivity, and $\Delta r$ is the thickness of the crust PRL_pagereddy . The calculation of $C_{V}$ and $\kappa$ is not straightforward, depending on pasta phases, superfluidity, magnetic field configuration, and the EOS PRL_pastabfieldcrustcooling2015 ; ChaikinKaminkerYakovlev2018 ; PotekhinChabrier2021 and is likely to be significantly modified in Scenarios B and C where the crust has more electrons and lacks free neutrons. Without making this detailed calculation, we rely on the quadratic dependence of the thermal relaxation timescale on the thickness of the crust and place the conservative constraint that the total crust thickness should be at least 0.2 km, roughly one fifth of its value in a normal neutron star.
@@ 6.1 Constraint from crust thermal relaxation in x-ray binaries
Scenario B gives the strongest constraint for small $\varepsilon$ from crust thermal relaxation because the pressure at the crust-core phase transition that determines at what $\varepsilon$ the crust appears is non-monotonic as a function of $K_{\sigma}$ (cf. Fig. 5). As $K_{\sigma}$ is increased from zero, the pressure for the crust-core phase transition decreases because the neutron matter surrounding nuclei in the inner crust becomes more attractive. At the same time, pressure in the outer crust that is dominated by electrons is increasing as nuclei become more neutron rich. The pressure at the crust-core phase transition will decrease as $K_{\sigma}$ is increased, giving a stronger bound on small $\varepsilon$, until the crust-core phase transition occurs at the same pressure as the transition from inner to outer crust (neutron drip pressure). The pressure at the crust-core phase transition will grow if $K_{\sigma}$ is increased further, giving a weaker bound on $\varepsilon$, since there are no more dripped neutrons in the crust and the pressure due to electrons increases with $K_{\sigma}$.
Figure 14: Constraint from crust thermal relaxation assuming a total crust thicknesses of $\Delta r<0.2$ km is disfavored by observation of thermal relaxation timescales. Three scenarios are shown, assuming no change to nuclear interactions (A), which we take as the conservative bound, and moderate (B) or large (C) increases of attractive nuclear interactions in $\theta=\pi$ matter (see Sec. 3 for details). The grey shaded regions are excluded by prior studies of the Sun and white dwarf stability Hook:2017psm ; Balkin:2022qer , black hole superradiance Baryakhtar:2020gao ; Hoof:2024quk and neutron star inspiral gravitational wave emission Zhang:2021mks . Recent constraints from isolated neutron star cooling nscooling_qcdaxion2024 are shown as a contour.
@@
6.2 Constraint from isolated neutron star cooling
The cooling of isolated neutron stars has been used to constrain light QCD axions in Ref. nscooling_qcdaxion2024 . Their constraint hinges on the envelope being thinner in a star with an axion-condensed crust than in a normal neutron star, providing less thermal insulation between the core and exterior of the star and shortening the cooling timescale of the core. Taking the boundary of the envelope to be at an energy density of $10^{10}$ g/cm3 when the energy density of the axion is not included, we find that for the range of $\varepsilon$ we consider, the entire envelope is in the domain wall region.
@@ 6.2 Constraint from isolated neutron star cooling
Figure 15 shows the thickness of the envelope as a function of $m_{a}$ and $\varepsilon$. In the white band, the phase transition to axion condensed matter occurs inside the star and the entire normal envelope is present. The feature around $\varepsilon\simeq 10^{-2}$ in all three panels is a result of a jump in the EOS just as the crust-core phase transition is at the bottom edge of the domain wall. Ref. nscooling_qcdaxion2024 found that when the envelope is at least a few times thinner than its normal $\sim 100$ m, the star cools too quickly and this region of parameter space is excluded. Taking a conservative estimate that the envelope must be at least 5 m thick to explain neutron star cooling data, we show the constraint for our three scenarios in Fig. 16. Notably, this indicates that the constraint in Ref. nscooling_qcdaxion2024 extends to lower axion masses and larger $\varepsilon$ than previously considered when the modification to the nuclear force is taken into account. This occurs because the chemical potential in the crust is significantly lower for larger $K_{\sigma}$ due to nuclei being more strongly bound. At a chemical potential so much lower than that of normal nuclei, the baryon density goes to zero much more rapidly as $\theta$ decreases in the domain wall, resulting in a thinner envelope. The envelope constraint applies for all $\varepsilon$ we consider where the domain wall is at the surface of the star (given by the first row of Table 2). At larger $\varepsilon$, the domain wall is inside the star and the entire normal envelope is present. Table 4 shows the constraint on $\varepsilon$ in our three scenarios for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$ where the domain wall becomes insignificant and the constraint is dependent on $\varepsilon$ only.
Figure 16: Constraint from isolated neutron star cooling based on the lack of an envelope in our three scenarios. We use the envelope thickness of $5$ m as a conservative minimum consistent with neutron star cooling. Our scenario (A) is in good agreement with the neutron star cooling bound of Ref. nscooling_qcdaxion2024 . Parameter space as in Fig. 14.
@@
6.3 Constraint from pulsar glitches
Pulsars are strongly magnetized, rapidly spinning neutron stars that are observed as a highly regular, periodic source of radiation. Some pulsars, however, have been observed to glitch, a phenomenon in which the pulse period rapidly diminishes on timescales of a day. This is typically understood as resulting from a rapid transfer of angular momentum from the core to the neutron superfluid in the crust of a normal neutron star andersonitoh_glitches1975 ; Ruderman_glitches1976 ; Pines_superfluidityinNS1985 . After including the effects of coupling the solid and superfluid components of the crust, this provides a lower limit on the fractional moment of inertia in the crust. The Vela pulsar provides the strongest constraint, requiring $I_{c}/I\gtrsim 7\%$ PRL_crustnotenough . This is a challenging constraint for many nuclear models to satisfy and it is unclear whether an additional reservoir of angular momentum may be required to explain the glitches of the Vela pulsar PRL_crustalentrainment_chamel ; PRC_crustmaybenough ; PRD_glitches_2016 . As such, although it is possible that the crust alone cannot account for the extreme glitch behavior of the Vela pulsar, we can conservatively exclude fractional moment of inertia less than 0.1%.
In our Scenarios B and C, there is no neutron superfluid when there is an axion-condensed crust since the crust-core phase transition occurs at a baryon density below the neutron drip. Since free neutrons dominate the crustal moment of inertia, these stars have a fractional moment of inertia that is an order of magnitude too small to explain the Vela pulsar glitches. In Scenario A, $\varepsilon\gtrsim 0.01$ is also excluded for the case of an axion condensed crust because the crust is mostly in the domain wall which is much thinner than a normal neutron star crust.
@@ 6.3 Constraint from pulsar glitches
Figure 17 shows the fractional of the moment of inertia of a $1.4\,M_{\odot}$ star as a function of $m_{a}$ and $\varepsilon$. For scenarios B and C, the moment of inertia in the crust is strongly suppressed when there is an axion-condensed crust (cf. Fig. 10). Again, the white band shows where the whole crust is present and there is no constraint, the hatched region shows where the domain wall approximation fails, and the solid black line divides the region where the domain wall is at the surface of the star and inside the star. The constraint that the fractional moment of inertia exceeds 0.1% is shown for our three scenarios in Fig. 18. Table 6 shows the constraint on $\varepsilon$ in our three scenarios for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$ where as before the domain wall does not play an important role and the constraint only depends on $\varepsilon$.
@@
6.4 $\mathbb{Z}_{\mathcal{N}}$ model
The discrete $\mathbb{Z}_{\mathcal{N}}$ symmetry proposed in Ref. Hook:2018jle provides a concrete realization of the exceptionally light QCD axion. This model predicts a suppression factor $\varepsilon$ of the axion mass for odd integer $\mathcal{N}$, given by
@@ 6.4 $\mathbb{Z}_{\mathcal{N}}$ model
where $z\equiv m_{u}/m_{d}$ and the large $\mathcal{N}$ limit has been taken DiLuzio:2021pxd . For small $\mathcal{N}$, the $\mathbb{Z}_{\mathcal{N}}$ model results in a modest change to the axion mass: for $\mathcal{N}=3$, $m_{a}/m_{a}^{\rm(QCD)}=0.94$. The first significant decrease comes at $\mathcal{N}=5$ giving $m_{a}/m_{a}^{\rm(QCD)}=0.65$ and $\mathcal{N}=11$ corresponds to approximately an order of magnitude suppression.
@@
In the $\mathbb{Z}_{\mathcal{N}}$ model for the exceptionally light QCD axion, the axion potential has a period of $2\pi/\mathcal{N}$ and the negative axion pressure is suppressed by an additional factor of $\mathcal{N}^{2}$ Hook:2018jle . When the domain wall is unimportant, the results are identical to our existing calculation but with a modified axion potential provided $\mathcal{N}$ is small enough that normal matter at saturation density is unstable to the nucleation of the $\theta=\pi$ phase. Thus, in the range $m_{a}\gtrsim 10^{-8}\,{\rm eV}$, the constraint on the $\mathbb{Z}_{\mathcal{N}}$ model parameter space can be inferred directly from the the constraint on $\varepsilon$ we have derived. The lower limit on $\mathcal{N}$ is determined from the difference in vacuum energy between the maximum at $\theta=\pi$ and minimum at $\theta=0$ of the potential given in Ref. DiLuzio:2021pxd , since the dependence on the relationship between $m_{a}$ and $f_{a}$ and the change to the period of the axion potential is minimal in this region. This lower limit on $\mathcal{N}$ for the various effects and scenarios is given in Table 7. For $m_{a}\gtrsim 10^{-8}\,{\rm eV}$, the combined constraint is the same for all three scenarios, $\mathcal{N}\geq 19$. Combined with the upper limit from white dwarf of $\mathcal{N}\leq 31$ Balkin:2022qer , this leaves a gap of $19\leq\mathcal{N}\leq 31$. On the other hand, the upper limit on $\mathcal{N}$ is also determined by the conditions for the appearance of the domain wall in Tab. 2. Although $\mathcal{N}=3$ and $5$ appear to be constrained in Scenario A and $\mathcal{N}=3$ appears to be constrained in Scenario B (see Table 7), for those choices of $\mathcal{N}$ and $K_{\sigma}$, the axion mass remains positive near $\theta=0$ in the $\mathbb{Z}_{\mathcal{N}}$ model. Therefore, even though the ground state of matter is axion condensed, since matter at saturation is only metastable in these cases, it is conceivable that the axion mass squared is always positive near $\theta=0$ anywhere in the neutron star, and these $\mathcal{N}$ cannot be conclusively excluded. However, it is possible that if axion condensation is favored at sufficiently large density where the axion mass squared turns negative near $\theta=0$, these $\mathcal{N}$ values could also be excluded in Scenarios A & B.
@@
For $m_{a}$ that is smaller than $10^{-8}\,{\rm eV}$, the distribution and property of the nuclear matter in the axion domain wall region is significantly affected by the change of the period of the axion potential in the $\mathbb{Z}_{\mathcal{N}}$ model, which requires further study. The parameter space in the gaps, as well as the lower axion mass region, might be probed by other considerations Hook:2019pbh of the formation of neutron stars during a supernova explosion. We leave these more detailed study within the $\mathbb{Z}_{\mathcal{N}}$ model to future work.
@@
Table 7: Minimum allowed $\mathcal{N}$ in the $\mathbb{Z}_{\mathcal{N}}$ model for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$. In Scenario A, the envelope constraint also permits $\mathcal{N}=3$. $\mathcal{N}$ must be an odd integer for the axion to solve the strong-CP problem in these models.

### v2

@@ V Finite size objects at zero pressure
An object with size $R\gg 1/m_{a}$ made of PNM is stable at almost zero pressure when
@@
$\begin{split}M&\simeq\frac{4\pi R^{3}}{3}\frac{m_{n}(5m_{n}\varepsilon m_{\pi}^{2}f_{\pi}^{2})^{3/5}}{(3\pi^{2})^{2/5}}\left[1-f(\pi)+\frac{m_{u}m_{d}}{2Rm_{a}(m_{u}+m_{d})^{2}}\right]^{3/5}\,,\\ &\simeq 2\times 10^{18}\varepsilon^{3/5}\left(\frac{R}{1\,\rm m}\right)^{3}\,{\rm kg}\,,\end{split}$ (34)
@@
Figure 11 shows the minimum stable radius for $\pi$-balls made of SNM and PNM and Fig. 12 shows the minimum stable mass, including nuclear interactions from the RMFT model and stabilized by axions. The horizontal dotted lines in Fig. 11 show when $m_{a}R=\pi$ and the domain wall approximation fails. The black dotted lines in Fig. 12 show the mass at which gravity cannot be neglected from Eq. (35). Note that when the constraints on axion parameter space from neutron star measurements we will derive in section VI are taken into account, these $\pi$-balls should typically have masses larger than the Earth. They will have masses comparable to heavy planets or brown dwarfs, and sizes as large as a few kilometers. Observation of one of these objects would be quite remarkable.
VI Observable implications and constraints
In this section, we estimate constraints on the exceptionally light axion parameter space from neutron star observations. We focus on three observables based on modifications to the crust of an axion-condensed neutron star: thermal relaxation times of the crust in accreting systems (Sec. VI.1), cooling rates of isolated neutron stars based on a diminished envelope, first considered in Ref. [40], (Sec. VI.2), and fractional moment inertia of the crust as inferred from pulsar glitches (Sec. VI.3). All of these rely only on the understanding of neutron star matter and dynamics below saturation density. In Sec. VI.4, we summarize the constraints in the context of the $Z_{\mathcal{N}}$ model of exceptionally light axions.
We note also that the mass-radius relation of axion condensed stars has some unusual features (see Sec. IV.1) which could lead to constraints on axion models. Current measurements of the masses and radii of neutron stars are not yet precise enough to draw conclusions. Next-generation experiments may significantly reduce the error bars and yield a useful constraint if the EOS for a range of densities, and for both $\theta=0$ and $\theta=\pi$, is better understood.
To understand the constraint from observation, we apply the methods discussed in Sec. IV and specifically the understanding of the axion domain wall established in Sec. IV.2. There are three distinct regions for this calculation. First, when $\theta=\pi$ is the favored phase all the way to zero pressure, the domain wall sits at the surface and forms a thin shell surrounding the rest of the star at negative bulk pressure, resulting in a thinner crust if $\varepsilon$ is not too small (cf. Fig. 10). Second, when the transition from $\theta=\pi$ to $\theta=0$ occurs in the crust of the normal phase, the domain wall connects a homogeneous axion condensed phase to a heterogeneous normal phase and the size of the crust is diminished. Third, when the phase transition to $\theta=\pi$ occurs at a density above the crust-core phase transition, the entire normal crust is present and we draw no constraint.
@@ VI Observable implications and constraints
DW in core at $n_{B}<n_{\rm sat}$ $0.380.27<\varepsilon<$ 0.38 $0.750.56<\varepsilon<$ 0.75 $10.96<\varepsilon<$ 1
VI.1 Constraint from crust thermal relaxation in x-ray binaries
Neutron stars in transiently accreting x-ray binaries go through periods of outburst and quiescence, reaching a quasi-steady state on astronomical timescales [83]. During outburst, matter is rapidly accreted onto the neutron star surface from a companion and pycnonuclear reactions deep in the crust heat the star. On timescales of tens to hundreds of days (see, e.g., Ref. [84]) the crust cools and comes back into thermal equilibrium with the core. Properties of the crust of neutron stars in these x-ray binaries can be inferred by observing the timescale for the crust to thermally relax following an outburst. The thermal timescale is roughly given by $\tau_{\rm th}\simeq C_{V}(\Delta r)^{2}/\kappa$ where $C_{V}$ is the specific heat of the crust, $\kappa$ is the thermal conductivity, and $\Delta r$ is the thickness of the crust [85]. The calculation of $C_{V}$ and $\kappa$ is not straightforward, depending on pasta phases, superfluidity, magnetic field configuration, and the EOS [86, 87, 88] and is likely to be significantly modified in Scenarios B and C where the crust has more electrons and lacks free neutrons. Without making this detailed calculation, we rely on the quadratic dependence of the thermal relaxation timescale on the thickness of the crust and place the conservative constraint that the total crust thickness should be at least 0.2 km, roughly one-fifth of its value in a normal neutron star.
@@ VI.1 Constraint from crust thermal relaxation in x-ray binaries
Scenario B gives the strongest constraint for small $\varepsilon$ from crust thermal relaxation because the pressure at the crust-core phase transition that determines at what $\varepsilon$ the crust appears is non-monotonic as a function of $K_{\sigma}$ (cf. Fig. 5). As $K_{\sigma}$ is increased from zero, the pressure for the crust-core phase transition decreases because the neutron matter surrounding nuclei in the inner crust becomes more attractive. At the same time, pressure in the outer crust that is dominated by electrons is increasing as nuclei become more neutron-rich. The pressure at the crust-core phase transition will decrease as $K_{\sigma}$ is increased, giving a stronger bound on small $\varepsilon$, until the crust-core phase transition occurs at the same pressure as the transition from inner to outer crust (neutron drip pressure). The pressure at the crust-core phase transition will grow if $K_{\sigma}$ is increased further, giving a weaker bound on $\varepsilon$, since there are no more dripped neutrons in the crust and the pressure due to electrons increases with $K_{\sigma}$.
Figure 14: Constraint from crust thermal relaxation assuming a total crust thickness of $\Delta r<0.2$ km is disfavored by observation of thermal relaxation timescales. Three scenarios are shown, assuming no change to nuclear interactions (A), which we take as the conservative bound, and moderate (B) or large (C) increases of attractive nuclear interactions in $\theta=\pi$ matter (see Sec. III for details). The grey shaded regions are excluded by prior studies of the Sun and white dwarf stability [36, 37], black hole superradiance [89, 90] and neutron star inspiral gravitational wave emission [39]. Recent constraints from isolated neutron star cooling [40] are shown as a contour. Plots of past constraints are generated with code from Ref. [10].
@@
VI.2 Constraint from isolated neutron star cooling
The cooling of isolated neutron stars has been used to constrain light QCD axions in Ref. [40]. Their constraint hinges on the envelope being thinner in a star with an axion-condensed crust than in a normal neutron star, providing less thermal insulation between the core and exterior of the star and shortening the cooling timescale of the core. Taking the boundary of the envelope to be at an energy density of $10^{10}$ g/cm3 when the energy density of the axion is not included, we find that for the range of $\varepsilon$ we consider, the entire envelope is in the domain wall region.
@@ VI.2 Constraint from isolated neutron star cooling
Figure 15 shows the thickness of the envelope as a function of $m_{a}$ and $\varepsilon$. In the white band, the phase transition to axion condensed matter occurs inside the star and the entire normal envelope is present. The feature around $\varepsilon\simeq 10^{-2}$ in all three panels is a result of a jump in the EOS just as the crust-core phase transition is at the bottom edge of the domain wall. Ref. [40] found that when the envelope is at least a few times thinner than its normal $\sim 100$ m, the star cools too quickly and this region of parameter space is excluded. Taking a conservative estimate that the envelope must be at least 5 m thick to explain neutron star cooling data, we show the constraint for our three scenarios in Fig. 16. Notably, this indicates that the constraint in Ref. [40] extends to lower axion masses and larger $\varepsilon$ than previously considered when the modification to the nuclear force is taken into account. This occurs because the chemical potential in the crust is significantly lower for larger $K_{\sigma}$ due to nuclei being more strongly bound. At a chemical potential much lower than that of normal nuclei, the baryon density goes to zero much more rapidly as $\theta$ decreases in the domain wall, resulting in a thinner envelope. The envelope constraint applies for all $\varepsilon$ we consider where the domain wall is at the surface of the star (given by the first row of Table 3). At larger $\varepsilon$, the domain wall is inside the star and the entire normal envelope is present. Table 5 shows the constraint on $\varepsilon$ in our three scenarios for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$ where the domain wall becomes insignificant and the constraint is dependent on $\varepsilon$ only.
Figure 16: Constraint from isolated neutron star cooling based on the lack of an envelope in our three scenarios. We use the envelope thickness of $5$ m as a conservative minimum consistent with neutron star cooling. Our scenario (A) is in good agreement with the neutron star cooling bound of Ref. [40]. Parameter space as in Fig. 14.
@@
VI.3 Constraint from pulsar glitches
Pulsars are strongly magnetized, rapidly spinning neutron stars that are observed as a highly regular, periodic source of radiation. Some pulsars, however, have been observed to glitch, a phenomenon in which the pulse period rapidly diminishes on timescales of a day. This is typically understood as resulting from a rapid transfer of angular momentum from the core to the neutron superfluid in the crust of a normal neutron star [91, 92, 93]. After including the effects of coupling the solid and superfluid components of the crust, this provides a lower limit on the fractional moment of inertia in the crust. The Vela pulsar provides the strongest constraint, requiring $I_{c}/I\gtrsim 7\%$ [94]. This is a challenging constraint for many nuclear models to satisfy and it is unclear whether an additional reservoir of angular momentum may be required to explain the glitches of the Vela pulsar [95, 96, 97]. As such, although it is possible that the crust alone cannot account for the extreme glitch behavior of the Vela pulsar, we can conservatively exclude fractional moment of inertia less than 0.1%.
In our Scenarios B and C, there is no neutron superfluid when there is an axion-condensed crust since the crust-core phase transition occurs at a baryon density below the neutron drip. Since free neutrons dominate the crustal moment of inertia, these stars have a fractional moment of inertia that is an order of magnitude too small to explain the Vela pulsar glitches. In Scenario A, $0.068>\varepsilon\gtrsim 0.01$ is also excluded for axion mass $m_{a}\gtrsim 10^{-9}\,\mbox{eV}$ because the crust is mostly in the domain wall which is much thinner than a normal neutron star crust.
@@ VI.3 Constraint from pulsar glitches
Figure 17 shows the fractional of the moment of inertia of a $1.4\,M_{\odot}$ star as a function of $m_{a}$ and $\varepsilon$. For scenarios B and C, the moment of inertia in the crust is strongly suppressed when there is an axion-condensed crust (cf. Fig. 10). Again, the white band shows where the whole normal crust is present and there is no constraint, the hatched region shows where the domain wall approximation fails, and the solid black line divides the region where the domain wall is at the surface of the star and inside the star. The constraint that the fractional moment of inertia exceeds 0.1% is shown for our three scenarios in Fig. 18. Table 7 shows the constraint on $\varepsilon$ in our three scenarios for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$, where, as before, the domain wall does not play an important role and the constraint only depends on $\varepsilon$.
@@
VI.4 $\mathbb{Z}_{\mathcal{N}}$ model
The discrete $\mathbb{Z}_{\mathcal{N}}$ symmetry proposed in Ref. [43] provides a concrete realization of the exceptionally light QCD axion. This model predicts a suppression factor $\varepsilon$ of the axion mass for odd integer $\mathcal{N}$, given by
@@ VI.4 $\mathbb{Z}_{\mathcal{N}}$ model
where $z\equiv m_{u}/m_{d}$ and the large $\mathcal{N}$ limit has been taken [44]. For small $\mathcal{N}$, the $\mathbb{Z}_{\mathcal{N}}$ model results in a modest change to the axion mass: for $\mathcal{N}=3$, $m_{a}/m_{a}^{\rm(QCD)}=0.94$. The first significant decrease comes at $\mathcal{N}=5$ giving $m_{a}/m_{a}^{\rm(QCD)}=0.65$ and $\mathcal{N}=11$ corresponds to approximately an order of magnitude suppression.
@@
In the $\mathbb{Z}_{\mathcal{N}}$ model for the exceptionally light QCD axion, the axion potential has a period of $2\pi/\mathcal{N}$ and the negative axion pressure is suppressed by an additional factor of $\mathcal{N}^{2}$ [43]. In the simple case where the domain wall connects $\theta=0$ and $\theta=\pi$ smoothly and the axion mass is sufficiently large that the domain wall pressure is unimportant, the results are identical to our existing calculation but with a modified axion potential provided $\mathcal{N}$ is small enough that normal matter at saturation density is unstable to the nucleation of the $\theta=\pi$ phase. Thus, in the range $m_{a}\gtrsim 10^{-8}\,{\rm eV}$, the constraint on the $\mathbb{Z}_{\mathcal{N}}$ model parameter space can be inferred directly from the constraint on $\varepsilon$ we have derived. The lower limit on $\mathcal{N}$ is determined from the difference in vacuum energy between the maximum at $\theta=\pi$ and minimum at $\theta=0$ of the potential given in Ref. [44] since the dependence on the relationship between $m_{a}$ and $f_{a}$ and the change to the period of the axion potential is minimal in this region. This lower limit on $\mathcal{N}$ for the various effects and scenarios is given in Table 8. For $m_{a}\gtrsim 10^{-8}\,{\rm eV}$, the combined constraint is the same for all three scenarios, $\mathcal{N}\geq 19$. Combined with the upper limit from white dwarf of $\mathcal{N}\leq 31$ [37], this leaves a gap of $19\leq\mathcal{N}\leq 31$. On the other hand, the upper limit on $\mathcal{N}$ is also determined by the conditions for the appearance of the domain wall in Tab. 3. Although $\mathcal{N}=3$ and $5$ appear to be constrained in Scenario A and $\mathcal{N}=3$ appears to be constrained in Scenario B (see Table 8), for those choices of $\mathcal{N}$ and $K_{\sigma}$, the axion mass remains positive near $\theta=0$ in the $\mathbb{Z}_{\mathcal{N}}$ model. Therefore, even though the ground state of matter is axion condensed, since matter at saturation is only metastable in these cases, it is conceivable that the axion mass squared is always positive near $\theta=0$ anywhere in the neutron star, and these $\mathcal{N}$ cannot be conclusively excluded. However, it is possible that if axion condensation is favored at a sufficiently large density where the axion mass squared turns negative near $\theta=0$, these $\mathcal{N}$ values could also be excluded in Scenarios A & B.
[…]
@@
For axion masses below $10^{-8}\,{\rm eV}$, the distribution and properties of nuclear matter in the axion domain wall region are significantly affected by the change of the period of the axion potential in the $\mathbb{Z}_{\mathcal{N}}$ model, which requires further study. The parameter space in the gaps, as well as the lower axion mass region, might be probed by other considerations of the formation of neutron stars during a supernova explosion [49]. We leave a more detailed study of the $\mathbb{Z}_{\mathcal{N}}$ model to future work.
@@
Table 8: Minimum allowed $\mathcal{N}$ in the $\mathbb{Z}_{\mathcal{N}}$ model for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$ when the domain wall connects $\theta=0$ and $\theta=\pi$. In Scenario A, the envelope constraint also permits $\mathcal{N}=3$. $\mathcal{N}$ must be an odd integer for the axion to solve the strong-CP problem in these models.